Q)The two blocks in an atwood machine have masses 2kg and 3 kg . find the work done by gravity during the fourth second after the system is released from rest?

Q)A block of mass 100 g is moved with a speed of 5 m/s at the highest point in a closed circular tube of radius 10 cm in a vertical plane. The cross- section of the tube is such that the block just fits in it. The block makes several oscillations inside the tube and finally stops at the lowest point. Find the work done by the tube on the block during the process.

Grade:12

1 Answers

anandpal shekhawat
14 years ago
Hi, A1. accn. of blocks will be g/5 = 2m/s^2 distance travelled by each block in 4th sec will be 7m work done by gravity on 3kg block = +3gh = 210J work done by gravity on 2kg block = -2gh = -140J so, net work done = 210+(-140) = +70J A2. Ei = (1/2)mu^2 + mgh = 1.35J Ef = 0 so, work done by tube = 0 - 1.35 = -1.35J thanks.